OBSERVATION OF THE Ã^{1}A′′ STATE OF ISOCYANOGEN

\underline{W. BRYAN LYNCH}, Department of Chemistry, University of Evansville, Evansville, IN 47722; HANS A. BECHTEL, ADAM H. STEEVES, JOHN J. CURLEY, and ROBERT W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139.

The Ã^{1}A′′ state of isocyanogen is observed using photofragment fluorescence excitation spectroscopy (PHOFEX). The spectra are highly congested, but progressions likely due to the Franck-Condon active C-N-C excited state bend are evident. Linewidth measurements indicate that the excited state lifetime is < 10 ps. These measurements are consistent with previous ab initio calculations, which predicted a bent excited state and a short lifetime due to predissociation. Although we do not believe the origin band is observed, we place an upper limit of 42523 cm^{−1} on the excited state energy.